# Graph-based Retrieval Augmented Generation for Dynamic Few-shot Text Classification

### v1

@@ 2.2.2. Data Augmentation-based models
$\mathbf{s}^{r}_{k}$ The score for all target labels at round $r$.
$y^{r*}_{k}$ The predicted label for the text.
$\mathcal{G}^{r}(\mathcal{V},\mathcal{E})$ Information graph with node set $\mathcal{V}$ and edge set $\mathcal{E}.$
@@
$\mathcal{D}^{r}$, $\mathcal{D}^{r}_{new}$ The labeled text for all labels and new labels at round $r$.
$\mathcal{T}^{r}$, $\mathcal{T}_{new}^{r}$ The text of all labels and new labels at round $r$.
$\mathcal{T}_{test}^{r}$ The testing text for at round $r$.
@@
$\mathcal{N}(\cdot)$ The neighbor set of a node in the information graph.
@@
Figure 2. An overview of the GORAG model is as follows. In Part 1, GORAG constructs a graph based on keywords extracted from the few-shot training data, taking into account the keyword extraction confidence and relatedness to the target labels. In Part 2, GORAG performs adaptive graph retrieval and utilizes the retrieved graph information to filter the target labels. In Part 3, GORAG carry out classification with the testing texts and the candidate labels retrieved from the information graph. In Part 4, GORAG applies online indexing to update the graph with newly extracted nodes from testing texts through online indexing.
2.2.3. LLM-based models
Recently, Large Language Model (LLM)-based models have undergone rapid development (Fernandez et al., 2023; Amer-Yahia et al., 2023; Zhang et al., 2024b; Miao et al., 2024) and have been successfully adapted to various tasks, including data discovery (Arora et al., 2023; Dong et al., 2023; Kayali et al., 2024), entity or schema matching (Zhang et al., 2023b; Tu et al., 2023; Fan et al., 2024; Du et al., 2024), and natural language to SQL conversion (Li et al., 2024; Ren et al., 2024; Trummer, 2022; Gu et al., 2023). Notably, LLMs are inherently capable of inference without fine-tuning (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024), making them originally suitable for dynamic text classification tasks. However, the lack of fine-tuning can lead LLMs to generate incorrect answers, as they lack task-specific knowledge, these incorrect answers are often referred to as hallucinations (Zhang et al., 2023c). To mitigate hallucinations, researchers have provided LLMs with side information for classification, namely long context RAG models. For example, Propositionizer (Chen et al., 2023) applies a fine-tuned LLM to convert side information into atomic expressions, to facilitate fine-grained information retrieval. RAPTOR (Sarthi et al., 2024) clusters side information and then generate summaries for each cluster, to help with the model prediction. However, the retrieved contents from these models remain unstructured and can be lengthy, which impedes the efficiency and effectiveness of LLMs, leading to lost-in-the-middle issue (Liu et al., 2024).
To address this, compression-based RAG models were proposed, these models try to compress side information to reduce the input length. Based on how they compress model inputs, these models can be classified into prompt compressor models, and graph-based RAG models. On the one hand, prompt compressor models, such as LLMLingua2 (Pan et al., 2024), apply LLM’s generation perplexity to filter out un-important tokens in the model input. Based on LLMLingua2, LongLLMLingua (Jiang et al., 2024) further considers the instruction to the model when compression, make the model instruction-aware. On the other hand, graph-based RAG models index side informations into a graph, and retrieve graph components or summaries, which are shorter than long contexts thet traditional RAG models retrieved. For instance, GraphRAG (Edge et al., 2024) constructs a graph and aggregates nodes into graph communities, then generate community summaries to help the LLM prediction. However, GraphRAG requires re-creating the community after each graph index, which incurs significant computational overhead.
To address this limitation of GraphRAG, two alternative models, HippoRAG (Gutiérrez et al., 2024) and LightRAG (Guo et al., 2024), have been proposed, which skip the formulation of communities and directly retrieve nodes or pathes from graphs. HippoRAG (Gutiérrez et al., 2024) retrieve nodes whose retrieval score is above a human defined threshold. However, the optimal retrieval threshold can vary across different data samples, and a globally fixed threshold selected by human may be suboptimal for the entire dataset. To address this issue, LightRAG (Guo et al., 2024) calculates the embedding similarity of text extracted entities with graph nodes, to achieve a one-to-one mapping and retrieve all triples involved these graph nodes.
3. Methdology
To address the aforementioned issues, we propose GORAG, a novel approach that achieves adaptive retrieval by extracting valuable side information from a minimum-cost spanning tree generated on the constructed graph. Specifically, GORAG takes into account each keyword’s with in the training and texting text corpus and relatedness to the class label by assigning different costs to edges on the constructed graph, allowing for more informed retrieval decisions. Moreover, GORAG incorporates an online indexing mechanism, which enables the model to index valuable information extracted from the testing texts to the constructed graph in real-time, thereby enhancing its ability to make accurate predictions in the future.
3.1. Framework Overview
As illustrated in Figure 2, our model consists of four primary components, i.e., Graph Construction, Graph Retrieval, Classification, and Online Indexing.
Figure 3. An example of the constructed information graph from WOS dataset, where red nodes denote the label node and black nodes denote the keyword node. In this subgraph, node Entity and node Neighborhood are considered most related to the label Relational Database, the edge between them with the label have the lowest cost assigned by our edge weighting mechanism.
Part 1: Graph Construction Graph construction proposes is to construct or update the information graph $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ at round $r$ based on the labeled data $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$ for at the $r$-th round. A small subgraph of the information graph created on WOS dataset is shown in Figure 3, graph formed like this will be used to provide retrieved-augmented information as context for unseen text, enabling LLMs to better understand the unseen text and accurately predict its labels. Specifically, under each new data $\mathcal{D}_{new}^{r}$ at the $r$-th round, we first extract the new keywords $\mathcal{K}_{new}^{r}\in\{t_{j}\}_{j=1}^{|\mathcal{D}_{new}^{r}|}$ from the few-shot training texts $\mathcal{T}^{r}_{new}=\{t_{j}\}_{j=1}^{|\mathcal{D}_{new}^{r}|}$. Then, we will then assign an edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ between each keyword $v\in\mathcal{K}_{new}^{r}$ and each label $y\in\mathcal{Y}^{r}_{new}$. We then compute the weight $w^{r}_{v,y}\in\mathcal{W}^{r}_{new}$ for each edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ based on the keyword’s importance with in the text corpus and relatedness to the label $y$. We then can merge all graphs $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ at each each round $r$ as the full graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$. More details can refer to Section 3.2.
Part 2: Graph Retrieval The Graph Retrieval process maps the extracted keyword nodes $\mathcal{V}^{t}$ from the unlabeled text $t$ to the constructed information graph $\mathcal{G}^{r}$ and generates a minimum-cost spanning tree on $\mathcal{G}^{r}$ that includes all these keywords. From this minimum-cost spanning tree, the candidate label set $\hat{\mathcal{Y}}_{t}^{r}$ is obtained, which is a reduced subset of the original target label set $\mathcal{Y}^{r}$. For more details, please refer to Section 3.3.
Part 3: Classification After retrieving the graph and obtaining the candidate types $\hat{\mathcal{Y}}_{t}^{r}$ for the text $t$ at round $r$, the classification procedure involves performing the final classification based on a large language model. This process utilizes the text $t$, the candidate labels $\hat{\mathcal{Y}}_{t}^{r}$ retrieved from the information graph $\mathcal{G}^{r}$, and descriptions $\mathcal{K}_{y_{i}}$ associated with each candidate label $y_{i}\in\hat{\mathcal{Y}}_{t}^{r}$. For more details, please refer to Section 3.4.1.
Part 4: Online Indexing The online indexing procedure dynamically indexes keywords, denoted as $\mathcal{V}_{\text{notexist}}^{t}$, which extracted from the testing text $t$ but are not in the existing graph $\mathcal{G}^{r}$. We then integrate these keywords into the graph to further enrich its structure, with weights assigned based on their importance within the text corpus and their relatedness to the predicted label $y^{*}_{t}$ for the testing text $t$. This process enhances the model’s ability to make more accurate predictions in the future. For more details, please refer to Section 3.4.2.
3.2. Part 1: Graph Construction
In this subsection, we would introduce the graph construction procedure of GORAG. To generate the information graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$ for each the $r$-th round, GORAG applies multiple instructions to the LLM for creating the graph node set $\mathcal{V}^{r}$ and the edges $e^{r}_{v,y}\in\mathcal{E}^{r}$ between keyword node $v$ and label node $y$. Also, we will assign weight $w^{r}_{v,y}\in\mathcal{W}^{r}$ for each each edge $e(v,y)$. The pseudo code of the graph construction algorithm is shown in Algorithm 1.
Algorithm 1 Graph Construction Algorithm of GORAG at the $r$-th round
@@ 3.2. Part 1: Graph Construction
0:  $\mathcal{D}_{new}^{r}$: Training text set at the $r$-th round and the provided label $y_{i}\in\mathcal{Y}_{new}^{r}$ for each training text $t_{j}$; $p_{extraction},\ p_{gen}$: The extraction and description generation instruction prompt; $\mathcal{G}^{r-1}$: Information graph of the previous round $r-1$.
0:  $\mathcal{G}^{r}$: Constructed information graph at the $r$-th round.
@@
At each $r$-th round, given the labeled training text $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, GORAG first extracts text keywords $\mathcal{K}^{r}_{new}$ from the text $\mathcal{T}^{r}_{new}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j}\}_{j=1}^{k}$ to serve as graph nodes. Specifically, we use the LLM model with an extraction instruction prompt $p_{extract}$, such as $p_{extract}=$Please extract some keywords from the following passage”. Then, we can get all text keywords $\mathcal{K}^{r}_{new}$ based on texts $\mathcal{T}^{r}_{new}$ as follows.
(2) $\mathcal{K}^{r}_{new}=\bigcup_{t\in\mathcal{T}^{r}_{new}}LLM(p_{extract},t).$
Also, by incorporating the new candidate labels $\mathcal{Y}^{r}_{new}$ at the $r$-th round, the graph node set $\mathcal{V}^{r}_{new}$ at the $r$-th round can be obtained as follows.
(3) $\mathcal{V}_{new}^{r}=\mathcal{K}_{new}^{r}\cup\mathcal{Y}_{new}^{r}.$
Then, GORAG links each keyword node $v\in\mathcal{K}_{{new}}^{r}$ to its corresponding label node $y\in\mathcal{Y}_{{new}}^{r}$, indicating that the keyword $v$ appears in texts associated with the label $y$.
We discuss how to compute the weight between each keyword node $v\in\mathcal{K}_{{new}}^{r}$ and each label $y\in\mathcal{Y}_{{new}}^{r}$. Considering the keywords $\mathcal{K}_{new}^{r}$ from the labeled text are not uniformly related to the text’s label $\mathcal{Y}_{new}^{r}$, we apply an edge weighting mechanism to assign a weight $w_{v,y}$ to each keyword-label link. This weight can be regarded as the correlation between keyword and each label. Firstly, we apply the $l_{2}$ normalized TF-IDF score (Sammut and Webb, 2010) to measure the importance and relatedness of a particular keyword $v\in\mathcal{K}_{new}^{r}$ and label $y\in\mathcal{Y}_{new}^{r}$ of the text $t\in\mathcal{T}^{r}$, where $\mathcal{T}^{r}=\cup_{i=1}^{r}\mathcal{T}_{new}^{i}$ denotes texts seen so far. Formally, given a keyword node $v\in\mathcal{K}_{new}^{r}$ and one text $t\in\mathcal{T}^{r}$, the correlation score $CS(v,t)$ between $v$ and $t$ is defined as follows.
(4) $CS(v,t)=\frac{count(v,t)}{|t|}\times log\frac{|\mathcal{T}^{r}|}{1+|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|},$
where $count(v,t)$ is the number of times that the term $v$ appears in the text $t$, and $|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|$ denotes the number of texts in the corpus $\mathcal{T}^{r}$ that contain the keyword $v$.
As the keyword $v$ can be extracted from multiple text source and from different rounds, the final edge weight $w_{v,y}^{r}$ of edge $e_{v,y}$ at the $r$-th round is calculated an the average of weights from all these texts. Formally, given the keyword $v\in\mathcal{K}_{new}^{r}$, label $y\in\mathcal{Y}_{new}^{r}$, and the texts $\mathcal{T}^{r,y}=\{t_{i}\}_{i=1}^{|\mathcal{T}^{r,y}|}$ consisting of texts $t_{i}$ associated with label $y$, the weight $w_{v,y}^{r}$ between $v$ and $y$ at the $r$-th round can be computed as follows:
(5) $w_{v,y}^{r}=\frac{\sum_{t_{j}\in\mathcal{T}^{r,y,v}}1-CS(v,t_{j})}{|\mathcal{T}^{r,y,v}|},$
where $\mathcal{T}^{r,y,v}=\{t_{j}|v\in t_{j}\land t_{j}\in\mathcal{T}^{r,y}\}$ is the text associated with label $y$ and containing the keyword $v$. We denote the generated graph for $\mathcal{K}_{new}^{r}$ and $\mathcal{Y}_{new}^{r}$ at the $r$-th round as $\mathcal{G}^{r}_{new}(\mathcal{V}_{new}^{r},\mathcal{E}_{new}^{r},\mathcal{W}_{new}^{r})$, where the node $\mathcal{V}^{r}_{new}=\mathcal{K}_{new}^{r}\cup\mathcal{Y}_{new}^{r}$ and edge $e^{r}_{v,y}\in\mathcal{E}_{new}^{r}$ is associated with the weight $w^{r}_{v,y}\in\mathcal{W}_{new}^{r}$.
We then discuss how to merge $\mathcal{G}^{r}_{new}$ into the graph from previous rounds $\mathcal{G}^{r-1}(\mathcal{V}^{r-1},\mathcal{E}^{r-1})$ to form the full graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r})$ as follows.
(6) $\mathcal{G}^{r}=Merge(\mathcal{G}^{r}_{new},\mathcal{G}^{r-1}),$
where $\mathcal{V}^{r}=\mathcal{V}^{r}_{new}\cup\mathcal{V}^{r-1}$, $\mathcal{E}^{r}=\mathcal{E}^{r}_{new}\cup\mathcal{E}^{r-1}$, and $\mathcal{W}^{r}=\mathcal{W}^{r}_{new}\cup\mathcal{W}^{r-1}$. Particularly, $\mathcal{G}^{0}=\emptyset$ and $\ r\geq 1$, Also, to guarantee the graph connectivity of the resulting graph, we add new edges between every newly added label node $y_{n}\in\mathcal{Y}_{new}^{r}$ and each old label node $y_{o}\in\mathcal{Y}^{r-1}$, the edge weight $w_{v,y}^{r}$ of edge $e_{y_{n},y_{o}}$ at round $r$ is calculated with the average weight of all edges that link keywords with label node $y_{n}$ or $y_{o}$, respectively:
(7) $w_{y_{n},y_{o}}^{r}=\frac{AvgWei(y_{n},\mathcal{G}^{r})+AvgWei(y_{o},\mathcal{G}^{r-1})}{2}$
(8) $AvgWei(y,\mathcal{G}^{r})=\frac{\sum_{v\in\mathcal{M}_{y}^{r}}w_{v,y}^{r}}{2\times|\mathcal{M}_{y}^{r}|},$
where $\mathcal{N}^{r}(y)$ denote the neighbor node set of label node $y$ in graph $\mathcal{G}^{r}$. Also, $\mathcal{M}_{y}^{r}=\{v\ |\ v\in\mathcal{N}^{r}(y)\ \land\ v\notin\mathcal{Y}^{r}\}$ denotes all neighbor nodes of label node $y$ that represent keywords, but not labels at the $r$-th round. After merge, the graph $\mathcal{G}^{r}$ would be used for future retrieval, and be further updated by GORAG’s online indexing mechanism later in Section 3.3.
@@
0:  $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$: The constructed information graph; $\mathcal{V}_{exist}^{t}\subseteq\mathcal{V}^{r}$: A set of keyword nodes extracted from text $t$ and can be mapped to graph $\mathcal{G}$; $\mathcal{Y}^{r}$: The target label set at the $r$-th round.

### v2

@@ Appendix A Notations
$\mathbf{s}^{r}$ The score for all target labels at round $r$.
$y^{r*}$ The predicted label for the text at round $r$.
$\mathcal{G}^{r}(\mathcal{V},\mathcal{E})$ weighted graph with node set $\mathcal{V}$ and edge set $\mathcal{E}.$
@@
$\mathcal{D}^{r}$, $\mathcal{D}^{r}_{new}$ The labeled training text for all seen labels at round $r$.
$\mathcal{T}^{r}$, $\mathcal{T}_{new}^{r}$ All seen texts at round $r$.
$\mathcal{T}_{test}^{r}$ The query text for at round $r$.
@@
$\mathcal{N}(\cdot)$ The neighbor set of a node in the weighted graph.
@@
Appendix B Complexity Analysis
In this section, we analysis the time and space complexity of GORAG’s graph construction, retrieval and prediction procedure. We denote the maximum number of terms of the input text as $m_{t}$, the number of unique terms of training corpus $\mathcal{D}_{new}^{r}$ as $u$, the LLM’s maximum input token length as $m_{i}$, LLM’s its maximum extraction, generation, and classification token length as $m_{e}$, $m_{g}$, and $m_{c}$, respectively.
B.1. Graph Construction Complexity
Firstly, we analysis the time and space complexity of GORAG’s graph construction. For the text keyword extraction procedure, the time complexity would be $O(|\mathcal{D}_{new}^{r}|(m_{i}+m_{e}))$; If label names are available, the time complexity of generat label descriptions and extract label keywords would be $O(|\mathcal{Y}^{r}_{new}|(m_{i}+m_{g}+m_{e}))$; Calculating the TFIDF and indexing edges to graph would require $O(|\mathcal{V}_{new}^{r}|(|\mathcal{D}_{new}^{r}|m_{t}+u))$ times, and merging graph requires
$O(|\mathcal{Y}^{r}_{new}||\mathcal{Y}^{r-1}|)$. Hence, the total time complexity of GORAG’s graph construction at the $r$-th round would be:
B.2. Retrieval and Classification Complexity
The time complexity of GORAG’s adaptive candidate type generation algorithm is the same with the Mehlhorn algorithm, which is $O(|\mathcal{E}^{r}|+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|)$ (Mehlhorn, 1988b); The time complexity of the online indexing mechanism would cost $|\mathcal{V}_{notexist}^{r}|(|\mathcal{T}_{test}^{r}|m_{t}+u)$; The time complexity of the final classification by LLM is $O(m_{i}+m_{c})$. Hence, the total time complexity of GORAG’s adaptive retrieval and classification can be denoted as
$O(|\mathcal{E}^{r}|+u|\mathcal{T}_{test}^{r}|).$
Appendix C Pseudo Code of GORAG
In this section, we provide the pseudo code of GORAG’s graph construction and adaptive candidate type generation algorithm in 1 and 2, respectively.
Appendix D Hardness Proof of The Adaptive Candidate Type Generation problem
Theorem D.1.
The Adaptive Candidate Type Generation problem is NP-hard.
Proof.
Algorithm 1 Graph Construction Algorithm of GORAG at round $r$
@@ Appendix D Hardness Proof of The Adaptive Candidate Type Generation problem
0:  $\mathcal{D}_{new}^{r}$: Training text set at the $r$-th round and the provided label $y_{i}\in\mathcal{Y}_{new}^{r}$ for each training text $t_{j}$; $p_{extraction},\ p_{gen}$: The extraction and description generation instruction prompt; $\mathcal{G}^{r-1}$: Weighted graph of the previous round $r-1$.
0:  $\mathcal{G}^{r}$: Constructed weighted graph at the $r$-th round.
@@
Appendix E Experiment Details
E.1. Further introduction of chosen baselines
NN-based Models
  • Entailment (Xia et al., 2021): Entailment concatenates the text data with each of the label names to form multiple entailment pairs with one text sample, hence increasing the number of training data and enhance its finetuning of a RoBERTa PLM (Liu, 2019). Entailment then carry out classification based on these entailment pairs and convert the text classification task into a binary classification task, for detecting whether the entailment pair formed is correct. When all entailment pairs of a particular text are considered to be incorrect, Entailment will directly give up the classification of this text. In this paper, we consider this classification as a hallucination.
Long Context RAG Models
Compression RAG Models
  • LongLLMLingua (Jiang et al., 2024a): LongLLMLingua is a instruction aware prompt compressor model, it applies LLM’s generation perplexity to filter out un-important tokens of the model input based on the retrieved side information and the task instruction.
  • GraphRAG (Edge et al., 2024): GraphRAG is a graph-Based RAG model that employs the LLM to extract texts’ entities and relations, which are then represented as nodes and edges in the weighted graph. GraphRAG then aggregates nodes into communities, and generates a comprehensive community report to encapsulate global information from texts.
@@ E.1. Further introduction of chosen baselines
0:  $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$: The constructed weighted graph; $\mathcal{V}_{exist}^{t}\subseteq\mathcal{V}^{r}$: A set of keyword nodes extracted from text $t$ and can be mapped to graph $\mathcal{G}$; $\mathcal{Y}^{r}$: The target label set at the $r$-th round.